OpenStudy (anonymous):

(30-a) / (2+4) has the value 3

mathslover (mathslover):

you mean to say : \[\large{\color{blue}{\frac{30-a}{2+4}=3}}\] \[\large{\color{red}{\textbf{We have to solve for a?}}}\]

OpenStudy (anonymous):

yes

OpenStudy (shubhamsrg):

2+4 ?? how can you solve that?? :O

mathslover (mathslover):

2+4=?

OpenStudy (anonymous):

6

OpenStudy (shubhamsrg):

and after that you multiply 6 on both sides!! omg,,a problem inside a problem! :(

OpenStudy (anonymous):

30-12/ (2+4) ,, how to solve that ?

mathslover (mathslover):

\[\large{\frac{30-a}{6}=3}\] \[\large{30-a=18}\] \[\large{a=30-18}\] what is 30-18?

OpenStudy (anonymous):

i got it, a is 12

mathslover (mathslover):

right and good

OpenStudy (anonymous):

and then how to solve 30-12/ (2+4)

OpenStudy (anonymous):

\[{30-12\over2+4}~~~ =3\]\[~~~~~~~~~~{18\over6}~ =3\]\[~~~~~~~~~~~~~~3=3\]

OpenStudy (anonymous):

ty i got it ;)

OpenStudy (anonymous):

youre welcome! :)

OpenStudy (anonymous):

30 - (12 / (2 + 4)) = 28

OpenStudy (anonymous):

36 / (a / 10) , i want answer less than 1 and greater than 36